Imagine a giant dry-cleaner's bag full of air at a temperature of -33 ∘C floating like a balloon with a string hanging from it 14 km above the ground.Estimate what its temperature would be if you were able to yank it suddenly back to Earth's surface.
1
Expert's answer
2019-12-06T09:42:39-0500

To return suddenly means that the process is adiabatic, i.e. "T_2 = T_1(\frac{p_2}{p_1})^{\frac{\gamma-1}{\gamma}} = 240(\frac{1}{0.15})^{0.29} =415 K."